A 50 ft rope is cut into two pieces. One piece is 2 ft longer than twice the length of the other. Find the length of the longer piece.
Answer by rm29924(97)

Let x = the shorter piece
the longer piece = 2x + 2
the two pieces added together = 50 ft
x + 2x + 2 = 50
3x + 2 = 50
3x = 48
x = 16
2x + 2
2(16) + 2
32 + 2 = 34
the longer piece is 34 ft